OpenStudy (anonymous):

7/2x-1=2x+5. solve the equation?

OpenStudy (anonymous):

have this instead... http://www.wolframalpha.com/input/?i=7%2F2x-1%3D2x%2B5

OpenStudy (anonymous):

note the Universe is R but {1/2} b/c that makes the denominator zero. now multiply both sides of the equation by 2x-1 you get 7=(2x+5)(2x-1) 7=4x2+8x−5 now subtract 7 from both sides you get 4x ^{2} + 8x -12=0 (which is a quadratic equation) now you can use the completing the square method or the quadratic formula to solve this you get (x-1)(x+3)=0 so x=1 and x=-3 are the solutions

OpenStudy (anonymous):

if you want to know how I solved the quadratic equation here is how, I have\[4x ^{2} + 8x -12=0\] I divided both sides by 4 I get \[x ^{2} + 2x -3=0\] now find two numbers when multiplied that give -3 and when added give 2 the numbers are 3 and -1 so the above quadratic equation can be written as \[x ^{2} + 3x -x -3\] factoring gives (x+3)(x-1)=0

OpenStudy (anonymous):

Done!

OpenStudy (mayankdevnani):

\[\frac{7}{2}x -1=2x+5\] \[\frac{7x-2}{2} =2x+5\] By Cross-Multiply_____ \[4x+10=7x-2\] \[-3x=-12\] \[x=4\]